### v1

@@ Abstract.
submissionid: 2923doi: XXXXXXX.XXXXXXXprice: 15.00isbn: XXX-X-XXXX-XXXX-X/XX/XXjournal: PACMHCIjournalyear: 2026journalvolume: 10journalnumber: 1article: GROUPxxpublicationmonth: 1copyright: noneccs: Human-centered computing Collaborative and social computing systems and toolsccs: Human-centered computing Interactive systems and toolsccs: Information systems Location based services
@@
[c]0.3 Refer to caption {subcaptionblock}[c]0.68 Refer to caption Refer to caption :
@@ 1. Introduction
Collective memory—originally a sociological construct—describes the active co-contribution, exchange, and visiting of community members’ individual memories (twymanBlackLivesMatter2017; harrisCollaborativeRecallCollective2008; wegnerTransactiveMemoryContemporary1987; jonesDesigningSocialMemory2016) that are related to the shared experiences of the community (sasDesigningCollectiveRemembering2006; ijabsCollectiveMemory2014). As an essential intangible possession of a community (harrisCollaborativeRecallCollective2008), collective memory differs from a simple collection of personal memories because it also emphasizes the interconnections among members’ shared experiences and life episodes, as well as how the interconnections benefit the community as a whole (sasDesigningCollectiveRemembering2006; harrisCollaborativeRecallCollective2008). On the one hand, it provides an ever-updating lens into the present conception of a group’s past, cultures, and vibes (roedigerCollectiveMemoryNew2015). On the other hand, it stimulates members to reflect, prospect, bond with each other (sasDesigningCollectiveRemembering2006), and eventually form group identities (harrisCollaborativeRecallCollective2008; eyermanPresentCultureTransmission2004; orianneCollectiveMemoryIndividual2023).
In the HCI and CSCW fields, we observe that most current work focuses on sociological analysis of existing communities’ collective memory (twymanBlackLivesMatter2017; ferronCollectiveMemoryBuilding2011; jonesCoconstructingFamilyMemory2018). They primarily see collective memory as a well-developed product through a holistic view of the entire community—what we refer to as an outcome-oriented, “top-down” perspective. What we think can generate complementary HCI insights is a process-oriented, “bottom-up” perspective of inquiry that prioritizes grassroots conceptualizations. For example, such inquiry can include community members’ own understanding of what constitutes collective memory, how collective memory inspires them, and how members practically co-contribute to collective memory in the everyday context. They shall provide more human-centered design insights for facilitating the community’s collective memory development rather than experts’ understandings. Therefore, this paper aims to practice this “bottom-up” perspective and uncover grassroots understanding and practices of this concept.
@@
In designing our research probes, we observed that the existing literature highlights collective memory development across two system modalities. One is the well-established content-oriented design of social platforms, such as online forums (glawionRememberingWorldWar2023). The other, while not so widely applied in the real world, emphasizes incorporating newer technologies, engaging memory-related interactions, and foregrounding interconnections among memories (velhinhoDesigningCollaborativeStorytelling2024; felasariCollatingCitysCollective2021). Therefore, we adopted a dual-system study setup to investigate whether different system designs influence people’s content generation practices and collective memory interpretations, and to ensure generalizability by accounting for potential preference variation. We designed a conventional content-oriented online forum system, as well as another locative and memory-skeuomorphic system derived from recent literature’s frameworks (liuGrassrootsHeritageCrisis2010; lakCollectiveMemoryUrban2019). And we deployed both systems as a between-subject study. Despite the dual system setup, we did not aim to find a better design from the two. Instead, we took an exploratory approach, using these probes to uncover people’s conceptualization of and interactions surrounding collective memory, as well as how different features might or might not support its development. Specifically, we are guided by the following research questions:
@@ 2.1. Collective Memory
Collective memory is originally a sociological construct describing shared conceptions of a community’s past and group members’ relations to it (heuxCollectiveMemoryAutobiographical2023). Although the precise definitions can vary with the disciplines and domains of analysis (harrisCollaborativeRecallCollective2008; heuxCollectiveMemoryAutobiographical2023), two essential elements in consensus are the interconnections among individual memories about the community and the dynamics of people interacting with each other’s memories to form a shared overall understanding of the community eventually (wegnerTransactiveMemoryContemporary1987). The first element—the interconnections—can be reflected by spatial or temporal similarities of memories (halbwachsSpaceCollectiveMemory1950), or that the memories are related to similar occurrences (sasDesigningCollectiveRemembering2006). For example, the collective memory of people in the same neighborhood is anchored by their similar experiences in the same physical surroundings, such as worship at the same church, trades at the same marketplace, etc. (halbwachsSpaceCollectiveMemory1950). Such shared context connects different personal narratives together, resulting in a combined micro- and macro-level overview of the community instead of a mere collection of narrative pieces (wegnerTransactiveMemoryContemporary1987). The second element—the dynamics—stresses the massive amount of visiting, exchanging, and complementing the shared knowledgebase, during which community members gain a deeper understanding of each other and the community as a whole (twymanBlackLivesMatter2017; harrisCollaborativeRecallCollective2008; wegnerTransactiveMemoryContemporary1987; jonesDesigningSocialMemory2016).
As briefed in Introduction, we categorize collective memory research into “top-down” and “bottom-up” perspectives. The former, being the mainstream adoption by sociologists, psychologists, and historians, holistically and retrospectively anatomizes a community with already-formed collective memory, often observing the group’s eventual remembrance in an outcome-oriented manner (e.g. (schwartzAbrahamLincolnForge2003; twymanBlackLivesMatter2017; halbwachsSpaceCollectiveMemory1950; wegnerTransactiveMemoryContemporary1987; ricoeurPersonalMemoryCollective2006)). Alternatively, the “bottom-up” and process-oriented perspective focuses on the grassroots practices of collective remembrance and understanding of collective memory as a shared possession (hirstCollectiveMemoryPsychological2018; heuxCollectiveMemoryAutobiographical2023). This perspective is less adopted in existing studies (heuxCollectiveMemoryAutobiographical2023). However, we would like to underline its potential for uncovering more first-person perceptions and opinions on collective memory co-construction, as well as echoing the emerging appeals for more human-centered designs that practically support the development of collective memory (jonesDesigningSocialMemory2016; vanhouseTechnologiesMemoryKey2008; liuGrassrootsHeritageCrisis2010).
@@ 2.2. HCI Approaches to Collective Memory
The former tends to stay in sync with sociologists’ and psychologists’ emphasis on the top-down investigation into existing collective memories (harrisCollaborativeRecallCollective2008; wegnerTransactiveMemoryContemporary1987; sasDesigningCollectiveRemembering2006; ijabsCollectiveMemory2014). For example, Ferron et al. and Twyman et al. conducted two studies on how communities collectively remember critical social movement events on Wikipedia (twymanBlackLivesMatter2017; ferronCollectiveMemoryBuilding2011). Two other studies investigated how collective memory was formed in families (jonesCoconstructingFamilyMemory2018) and trauma teams (sarcevicTransactiveMemoryTrauma2008). They have highlighted the importance of digitalizing and preserving personal narratives and fostered the interactions among community members, but have discussed less how to practically tackle the fragmented nature of memory pieces. Also, they focused less on communities’ daily routines and ordinary experiences, another organic part of the collective memory (halbwachsSpaceCollectiveMemory1950).
@@
In comparison, Jones et al. propose to approach collective memory as a procedure of continuous “(re)construction and (re)interpretation”, especially the use of artifacts to facilitate “creating, sharing, and revisiting” of the collective memory (jonesDesigningSocialMemory2016). Accordingly, relevant designs strove to enable effective memory aggregation and retrieval within communities (hakkilaConnectingEvaluatingIndigenous2023; klaebeDigitalStorytellingHistory2007; lindleyNarrativeMemoryPractice2009), as well as integrated memory authoring and exploration flow (wegnerTransactiveMemoryContemporary1987). Uriu et al. fostered serendipitous discoveries of memories in the extended family tree (uriuCaraClockInteractivePhoto2009). Although it targeted small-scale kinship interactions, it has demonstrated the value of serendipity and assisted roaming over the collectivity. Cheverst et al. showed the advantages of locative media experience for collective memory with a locative authoring and exploration application for a rural village community (cheverstSupportingConsumptionCoauthoring2017). Being situated in the physical surroundings created a more immersive and empathetic experience of collective memory exploration (cheverstSupportingConsumptionCoauthoring2017). Nonetheless, their work advocated an expert-directed curation of collective memory, and their evaluation mostly covers usability topics (cheverstSupportingConsumptionCoauthoring2017). Another system by Cranshaw et al. persuading travelers to journal for less prominent places in a travel destination (cranshawJourneysNotesDesigning2016). Though they provided exemplary designs of an interactive and intentionally poetic system to engage the community, their evaluation did not cover how the design actually influenced the journaling experience.
@@
In sum, community members’ own perception of collective memory is less discussed in current works, nor its relationships with the community’s interactions (halbwachsSpaceCollectiveMemory1950). This is exactly the focus of this paper, and we hope to ultimately reinforce the bridge between the theoretical framing of collective memory and the grassroots efforts in its active development.
@@ 2.3. Individual Narratives of the Past
There are other HCI works that support the recording and revisiting of individual memories, using forms such as tiles (bennettTopoTilesStorytellingCare2015), buttons (riggsDesigningArchiveFeelings2024), and paintings (wanMetamorpheusInteractiveAffective2024) as symbolic representations and proxies of personal memories, and digital replications of past experiences in AR (liExploringOpportunitiesAR2023; merrimanFamiliarEnvironmentsEnhance2016). Their main focus was the outcome itself as individual memorial artifacts that are circulated within a restricted scope. They did not emphasize the dynamic, transactive process of a community’s memories and the formation of an integrated conception. Still, these studies provided implications on how to stimulate memory with situated environments (bennettTopoTilesStorytellingCare2015), photos (liExploringOpportunitiesAR2023), and memory-related objects (riggsDesigningArchiveFeelings2024).
Meanwhile, we note that social media sites are also common places to publish and archive personal experiences, as well as to disseminate them within the shared social network. However, they are still designed to highlight people’s individual lives at the moment (louieOpportunisticCollectiveExperiences2021) with limited advocacy for memory contents that are meaningful to a community or for interpreting contents from the community’s perspective (naamanItReallyMe2010).
@@ 3. Prototyping Mobile Apps for Collective Memory Co-Contribution
As mentioned in section 1, considering the transactive and social nature of collective memory (wegnerTransactiveMemoryContemporary1987), we employed a dual-system setup to both respect common existing practices and explore recent design proposals and frameworks. We adopted both a conventional forum-like system and a dedicatedly designed system (namely Drop’n’Collect), in the hope of better probing into users’ interpretations of collective memory, practices relevant to its co-contribution, and design opportunities to facilitate the process. The next two subsections cover our design process for Drop’n’Collect. While the forum system did not receive a dedicated design process (because it aligns with the common forum user experience), we ensured that both systems’ features paired up to serve the same set of user intents. The final features of both systems are described pairwise in the last subsection.
@@ 3.1. Design Goals and Requirements for Drop’n’Collect
Referring to Liu’s framework (liuGrassrootsHeritageCrisis2010), we highlighted the following five design goals for HCI systems to support “socially-distributed curation” of collective memories: (1) active participation at the individual level, (2) content collection at the community level, (3) continuous active efforts across time, (4) meaning-making of the interconnected memories, and (5) interactive memory sharing.
@@
We first translated these high-level goals into actionable design requirements (Figure 4). To start with, community members’ active contribution (Goal 1) is intrinsically motivated upon recollecting their community-related experiences and feelings (cranshawJourneysNotesDesigning2016). As the recollected episodes and emotions may disappear when interrupted (viskontasNeuralCorrelatesRecollection2009), it is important to nudge people to put them down when freshly in the mood. We thus summarized the design requirement: DR1. evoke memories in members and prompt recording in-situ. Second, a more complete picture of the group can be obtained by involving more people to contribute (Goal 2), as every member’s memory only reflects certain facets of the community (wegnerTransactiveMemoryContemporary1987; heuxCollectiveMemoryAutobiographical2023). In this regard, we proposed: DR2. ensure coverage and diversity of memories at a community scale. Third, a living collective memory requires continuous contributions of active members (Goal 3). If the task is very demanding or boring, it is difficult to maintain prolonged involvement. This led to DR3. retain participation by balancing the load and the necessary memory-recording requirements. Fourth, to ensure that individual memories are interconnected in the community, it is critical to encourage community members to seek meanings from such connections and consciously select individual memories (Goal 4). The most intuitive way to achieve it is by situating an individual’s memory in the community’s unique context. We framed it as DR4. associating memories through their common dimensions (e.g., locations, activities). Fifth, from the user experience perspective, designing an interactive story-sharing experience (Goal 5) could enhance listeners’ engagement and emotional impact (pittarelloDesigningContextAwareArchitecture2011). We kept it straightforward as DR5. interactive story sharing. A design requirement can also meet multiple goals. For example, ensuring memory coverage (DR2) may lay a foundation for meaning-making of memories’ interconnections (Goal 4); easing the system’s load (DR3) may invite more concurrent participation as well (Goal 2); connecting members memories (DR4) may help members see the value of their contributions and consequently encourage their participation (Goal 1) and retention (Goal 3). Interactive sharing (DR5) can become engaging enough to motivate individuals’ and the community’s maintained participation (Goals 1 & 3).
Figure 4. An overview of the design process for Drop’n’Collect.:

### v2

@@ Abstract.
copyright: nonejournal: PACMHCIjournalyear: 2026journalvolume: 10journalnumber: 1article: GROUP009publicationmonth: 3doi: 10.1145/3799444submissionid: 2923ccs: Human-centered computing Collaborative and social computing systems and toolsccs: Human-centered computing Interactive systems and toolsccs: Information systems Location based services
@@
[c]0.3 Refer to caption {subcaptionblock}[c]0.68 Refer to caption Refer to caption (a) The screenshot of Drop’n’Collect home page. A map view shows the current location at the center and three location pins to indicate nearby posts. A horizontal scrolling list displays three memory postcards in sepia filter. Some orange circled numbers on the map show how many posts are in adjacent areas. (b) Two photos. The first is a user holding a mobile phone in front of a yellow building on the road. The mobile phone shows a memory postcard, which is exactly the same spot but the building and the road are destroyed by natural disasters. The second is a screenshot of the system showing the same memory post. It shows the flip side of the memory postcard, which is a user’s memory about a typhoon.
@@ 1. Introduction
Collective memory—originally a sociological construct—describes the active co-contribution, exchange, and visiting of community members’ individual memories (Twyman et al., 2017; Harris et al., 2008; Wegner, 1987; Jones, 2016) that are related to the shared experiences of the community (Sas and Dix, 2006; Ijabs, 2014). As an essential intangible possession of a community (Harris et al., 2008), collective memory differs from a simple collection of personal memories because it also emphasizes the interconnections among members’ shared experiences and life episodes, as well as how the interconnections benefit the community as a whole (Sas and Dix, 2006; Harris et al., 2008). On the one hand, it provides an ever-updating lens into the present conception of a group’s past, cultures, and vibes (Roediger and Abel, 2015). On the other hand, it stimulates members to reflect, prospect, bond with each other (Sas and Dix, 2006), and eventually form group identities (Harris et al., 2008; Eyerman, 2004; Orianne and Eustache, 2023).
In the HCI and CSCW fields, we observe that most current work focuses on sociological analysis of existing communities’ collective memory (Twyman et al., 2017; Ferron and Massa, 2011; Jones and Ackerman, 2018). They primarily see collective memory as a well-developed product through a holistic view of the entire community—what we refer to as an outcome-oriented, “top-down” perspective. What we think can generate complementary HCI insights is a process-oriented, “bottom-up” perspective of inquiry that prioritizes grassroots conceptualizations. For example, such inquiry can include community members’ own understanding of what constitutes collective memory, how collective memory inspires them, and how members practically co-contribute to collective memory in the everyday context. They shall provide more human-centered design insights for facilitating the community’s collective memory development rather than experts’ understandings. Therefore, this paper aims to practice this “bottom-up” perspective and uncover grassroots understanding and practices of this concept.
@@
In designing our research probes, we observed that the existing literature highlights collective memory development across two system modalities. One is the well-established content-oriented design of social platforms, such as online forums (Glawion, 2023). The other, while not so widely applied in the real world, emphasizes incorporating newer technologies, engaging memory-related interactions, and foregrounding interconnections among memories (Velhinho et al., 2024; Felasari and Peng, 2021). Therefore, we adopted a dual-system study setup to investigate whether different system designs influence people’s content generation practices and collective memory interpretations, and to ensure generalizability by accounting for potential preference variation. We designed a conventional content-oriented online forum system, as well as another locative and memory-skeuomorphic system derived from recent literature’s frameworks (Liu, 2010; Lak and Hakimian, 2019). And we deployed both systems as a between-subject study. Despite the dual system setup, we did not aim to find a better design from the two. Instead, we took an exploratory approach, using these probes to uncover people’s conceptualization of and interactions surrounding collective memory, as well as how different features might or might not support its development. Specifically, we are guided by the following research questions:
@@ 2.1. Collective Memory
Collective memory is originally a sociological construct describing shared conceptions of a community’s past and group members’ relations to it (Heux et al., 2023). Although the precise definitions can vary with the disciplines and domains of analysis (Harris et al., 2008; Heux et al., 2023), two essential elements in consensus are the interconnections among individual memories about the community and the dynamics of people interacting with each other’s memories to form a shared overall understanding of the community eventually (Wegner, 1987). The first element—the interconnections—can be reflected by spatial or temporal similarities of memories (Halbwachs, 1950), or that the memories are related to similar occurrences (Sas and Dix, 2006). For example, the collective memory of people in the same neighborhood is anchored by their similar experiences in the same physical surroundings, such as worship at the same church, trades at the same marketplace, etc. (Halbwachs, 1950). Such shared context connects different personal narratives together, resulting in a combined micro- and macro-level overview of the community instead of a mere collection of narrative pieces (Wegner, 1987). The second element—the dynamics—stresses the massive amount of visiting, exchanging, and complementing the shared knowledgebase, during which community members gain a deeper understanding of each other and the community as a whole (Twyman et al., 2017; Harris et al., 2008; Wegner, 1987; Jones, 2016).
As briefed in Introduction, we categorize collective memory research into “top-down” and “bottom-up” perspectives. The former, being the mainstream adoption by sociologists, psychologists, and historians, holistically and retrospectively anatomizes a community with already-formed collective memory, often observing the group’s eventual remembrance in an outcome-oriented manner (e.g. (Schwartz, 2003; Twyman et al., 2017; Halbwachs, 1950; Wegner, 1987; Ricoeur, 2006)). Alternatively, the “bottom-up” and process-oriented perspective focuses on the grassroots practices of collective remembrance and understanding of collective memory as a shared possession (Hirst et al., 2018; Heux et al., 2023). This perspective is less adopted in existing studies (Heux et al., 2023). However, we would like to underline its potential for uncovering more first-person perceptions and opinions on collective memory co-construction, as well as echoing the emerging appeals for more human-centered designs that practically support the development of collective memory (Jones, 2016; Van House and Churchill, 2008; Liu, 2010).
@@ 2.2. HCI Approaches to Collective Memory
The former tends to stay in sync with sociologists’ and psychologists’ emphasis on the top-down investigation into existing collective memories (Harris et al., 2008; Wegner, 1987; Sas and Dix, 2006; Ijabs, 2014). For example, Ferron et al. and Twyman et al. conducted two studies on how communities collectively remember critical social movement events on Wikipedia (Twyman et al., 2017; Ferron and Massa, 2011). Two other studies investigated how collective memory was formed in families (Jones and Ackerman, 2018) and trauma teams (Sarcevic et al., 2008). They have highlighted the importance of digitalizing and preserving personal narratives and fostered the interactions among community members, but have discussed less how to practically tackle the fragmented nature of memory pieces. Also, they focused less on communities’ daily routines and ordinary experiences, another organic part of the collective memory (Halbwachs, 1950).
@@
In comparison, Jones et al. propose to approach collective memory as a procedure of continuous “(re)construction and (re)interpretation”, especially the use of artifacts to facilitate “creating, sharing, and revisiting” of the collective memory (Jones, 2016). Accordingly, relevant designs strove to enable effective memory aggregation and retrieval within communities (Häkkilä et al., 2023; Klaebe et al., 2007; Lindley et al., 2009), as well as integrated memory authoring and exploration flow (Wegner, 1987). Uriu et al. fostered serendipitous discoveries of memories in the extended family tree (Uriu et al., 2009). Although it targeted small-scale kinship interactions, it has demonstrated the value of serendipity and assisted roaming over the collectivity. Cheverst et al. showed the advantages of locative media experience for collective memory with a locative authoring and exploration application for a rural village community (Cheverst et al., 2017). Being situated in the physical surroundings created a more immersive and empathetic experience of collective memory exploration (Cheverst et al., 2017). Nonetheless, their work advocated an expert-directed curation of collective memory, and their evaluation mostly covers usability topics (Cheverst et al., 2017). Another system by Cranshaw et al. persuading travelers to journal for less prominent places in a travel destination (Cranshaw et al., 2016). Though they provided exemplary designs of an interactive and intentionally poetic system to engage the community, their evaluation did not cover how the design actually influenced the journaling experience.
@@
In sum, community members’ own perception of collective memory is less discussed in current works, nor its relationships with the community’s interactions (Halbwachs, 1950). This is exactly the focus of this paper, and we hope to ultimately reinforce the bridge between the theoretical framing of collective memory and the grassroots efforts in its active development.
@@ 2.3. Individual Narratives of the Past
There are other HCI works that support the recording and revisiting of individual memories, using forms such as tiles (Bennett et al., 2015), buttons (Riggs et al., 2024), and paintings (Wan et al., 2024) as symbolic representations and proxies of personal memories, and digital replications of past experiences in AR (Li et al., 2023; Merriman et al., 2016). Their main focus was the outcome itself as individual memorial artifacts that are circulated within a restricted scope. They did not emphasize the dynamic, transactive process of a community’s memories and the formation of an integrated conception. Still, these studies provided implications on how to stimulate memory with situated environments (Bennett et al., 2015), photos (Li et al., 2023), and memory-related objects (Riggs et al., 2024).
Meanwhile, we note that social media sites are also common places to publish and archive personal experiences, as well as to disseminate them within the shared social network. However, they are still designed to highlight people’s individual lives at the moment (Louie et al., 2021) with limited advocacy for memory contents that are meaningful to a community or for interpreting contents from the community’s perspective (Naaman et al., 2010).
@@ 3. Prototyping Mobile Apps for Collective Memory Co-Contribution
As mentioned in section 1, considering the transactive and social nature of collective memory (Wegner, 1987), we employed a dual-system setup to both respect common existing practices and explore recent design proposals and frameworks. We adopted both a conventional forum-like system and a dedicatedly designed system (namely Drop’n’Collect), in the hope of better probing into users’ interpretations of collective memory, practices relevant to its co-contribution, and design opportunities to facilitate the process. The next two subsections cover our design process for Drop’n’Collect. While the forum system did not receive a dedicated design process (because it aligns with the common forum user experience), we ensured that both systems’ features paired up to serve the same set of user intents. The final features of both systems are described pairwise in the last subsection.
@@ 3.1. Design Goals and Requirements for Drop’n’Collect
Referring to Liu’s framework (Liu, 2010), we highlighted the following five design goals for HCI systems to support “socially-distributed curation” of collective memories: (1) active participation at the individual level, (2) content collection at the community level, (3) continuous active efforts across time, (4) meaning-making of the interconnected memories, and (5) interactive memory sharing.
@@
We first translated these high-level goals into actionable design requirements (Figure 4). To start with, community members’ active contribution (Goal 1) is intrinsically motivated upon recollecting their community-related experiences and feelings (Cranshaw et al., 2016). As the recollected episodes and emotions may disappear when interrupted (Viskontas et al., 2009), it is important to nudge people to put them down when freshly in the mood. We thus summarized the design requirement: DR1. evoke memories in members and prompt recording in-situ. Second, a more complete picture of the group can be obtained by involving more people to contribute (Goal 2), as every member’s memory only reflects certain facets of the community (Wegner, 1987; Heux et al., 2023). In this regard, we proposed: DR2. ensure coverage and diversity of memories at a community scale. Third, a living collective memory requires continuous contributions of active members (Goal 3). If the task is very demanding or boring, it is difficult to maintain prolonged involvement. This led to DR3. retain participation by balancing the load and the necessary memory-recording requirements. Fourth, to ensure that individual memories are interconnected in the community, it is critical to encourage community members to seek meanings from such connections and consciously select individual memories (Goal 4). The most intuitive way to achieve it is by situating an individual’s memory in the community’s unique context. We framed it as DR4. associating memories through their common dimensions (e.g., locations, activities). Fifth, from the user experience perspective, designing an interactive story-sharing experience (Goal 5) could enhance listeners’ engagement and emotional impact (Pittarello, 2011). We kept it straightforward as DR5. interactive story sharing. A design requirement can also meet multiple goals. For example, ensuring memory coverage (DR2) may lay a foundation for meaning-making of memories’ interconnections (Goal 4); easing the system’s load (DR3) may invite more concurrent participation as well (Goal 2); connecting members memories (DR4) may help members see the value of their contributions and consequently encourage their participation (Goal 1) and retention (Goal 3). Interactive sharing (DR5) can become engaging enough to motivate individuals’ and the community’s maintained participation (Goals 1 & 3).
Refer to caption A diagram with four columns: ”Design Goals”, ”Design Requirements”, ”Design Approaches”, and ”Features”. The items and the linkage between them are stated in the body text.
Figure 4. An overview of the design process for Drop’n’Collect.